# Aragog: Just-in-Time Model Routing for Scalable Serving of Agentic Workflows

### v1

@@ 4 Implementation
Figure 10: Aragog’s DSPy program wrapper abstraction for enabling graph structure extraction and stage-wise scheduling.
We implement Aragog as an orchestration layer in 6.8K lines of Python and Triton code that sits between agentic workflow front end and LLM serving engines without the need of modifying both.
Frontend and Workflow Specification. As shown in Figure 10, Aragog converts any DSPy program into its internal workflow representation, enabling runtime model reconfiguration and asynchronous stage-wise execution. We choose DSPy for its lightweight design and ease of use, making it popular for building agentic workflows. However. Aragog can support other agentic frameworks with a specific implementation to extract the workflow graph.
Router Inference We implement dynamic batching for shared embedding model serving. To accelerate classifier inference, we develop custom Triton kernels that fuse classifiers across a tunable number of classifiers into a single batched operation. We leverage CUDA graphs to further reduce the kernel launching overheads.
Runtime Scheduler. The scheduler runs asynchronously. It maintains: (1) FIFO request queues with look-ahead scheduling, (2) per-model engine states tracking batch occupancy, (3) configuration caches mapping requests to viable sets, and (4) exponential moving averages of queue metrics. We implement beam search using priority queues, where each state encodes partial assignments as bit vectors for efficient manipulation.
Serving Backend Integration. Aragog use SGLang [sglang], a widely used serving framework for large language models as our serving backend. However, Aragog naturally supports any serving backends that provide OpenAI-compatible APIs (vLLM [vllm], TensorRT-LLM [tensorrt-llm] and HuggingFace TGI [TGI]).
@@ 5 Evaluation
We evaluated Aragog across a wide range of agentic workloads and varying request rates. Our key findings are:
@@
    Aragog improves maximum serving throughput by 50.0–76.3% and 78.1–217.0% over per-input and per-workflow optimizations, respectively.
@@
    Aragog maintains lower latency under increasing load. Under the highest load for all workloads, Aragog reduces P25, median, and P95 end-to-end latency by 23.6–54.6%, 32.5–58.8%, and 46.8–66.1% over per-input optimization and 58.4–75.1%, 60.0–78.9%, and 63.2–81.5% over per-workflow optimization, respectively.
@@
    Aragog preserves accuracy compared to always using the largest model, with maximum degradation of 1%.
@@
    Aragog demonstrates robust performance improvements over baselines across diverse model families, router architectures, configuration space sizes, and GPU allocation.
@@ 5.1 Experimental Setup
Models. We use the Qwen-2.5 model family [qwen]: 7B, 14B, 32B parameters. This provides a diverse range of capability-cost tradeoffs for each agent in the workflows. We use Snowflake Arctic-Embed-L-v2.0 [snowflake-model] as the router’s embedding model, which has a small number of parameters while providing high-quality embeddings and supporting relatively long context lengths. Additionally, we explore more diverse model family setups and different embedding models in §5.3.
Testbed. We conduct our experiments on an H100 GPU node, which contains 8 NVIDIA H100 80GB GPUs interconnected via NVLink and a 96-core Intel Sapphire Rapids CPU. We use SGLang as our model serving engine. For models larger than 7B parameters, we enable tensor parallelism with a degree of 2 for 14B models and 4 for 32B models, following standard practice. We dedicate the one remained GPU to router inference. For baselines without routing components, this GPU instead serves an additional 7B model to ensure fair resource allocation. We also explore different initial GPU allocation configurations in §5.3. As in prior work [vllm, sglang, autellix], we simulate request arrivals using a Poisson process with varying arrival rates to capture different load conditions.
Workflows and datasets. We evaluate Aragog across diverse workflow patterns and application domains.
@@
    Self-refine for coding on HumanEval [humaneval] (164 programming problems), using a three-agent workflow with generation, critique, and refinement stages.
@@
    Text-to-SQL on Bird-SQL [bird] (780 samples), where query construction is decomposed into keyword extraction, column selection, SQL generation, and refinement stages.
@@
    Majority voting [self-consistency] (four parallel voters and one aggregator) on MMLU-Pro [mmlu-pro] (1,500 sampled questions), where multiple reasoning paths are generated in parallel and aggregated.
@@
    Task decomposition on StrategyQA [strategyQA] (1,500 sampled questions), where problems first pass through decomposer, then four sub-task solver in parallel, and finally answer aggregator.
Baselines. We compare Aragog against two oracle baselines that represent optimal performance for each category: (1) Per-workflow optimization: selects the cheapest configuration among all configurations that match the most expensive configuration’s accuracy, the configuration stays fixed for all requests; and (2) Per-input optimization: performs router-based selection for each individual request. Both baselines are given an oracle with perfect accuracy knowledge for serving requests, representing the performance upper bound of each approach. Additionally, we enforce the same router inference overhead for the per-input baseline as Aragog. Note that this routing overhead is already underestimated in favor of the per-input baseline’s single monolithic router, as discussed in §3.1. Moreover, all baselines incorporate the same state-of-the-art graph-aware scheduling for agentic workflows as Aragog (§3.2).
Figure 11: Average latency and throughput comparison under varying request rates and across five workflow dataset pairs. Each dot represents the average latency and throughput under a specific request rate.
Figure 12: Median latencies for different workflow dataset pairs under varying request rates, with error bars for P25 and P95 latencies
@@ 5.2 End-to-End Performance
Figures 1113 compare Aragog with two baselines that represent optimal performance for each category across diverse workloads. Overall, Aragog significantly improves the maximal serving capacity and lowers latencies, while matching the accuracy when always using the largest model.
@@
Serving capacity improvement. In Figure 11, we evaluate the end-to-end serving capacity of Aragog against baselines across five workloads. For each workload, we sweep the incoming request rate from low to high and measure the achieved throughput and average latency at each rate, shown as individual points on the curves. Aragog consistently achieves the highest throughput and lowest average latency under varying request rates. Specifically, Aragog improves maximum serving throughput by 50.0–76.3% over per-input optimization and 78.1–217.0% over per-workflow optimization. Additionally, Aragog reduces average latency under the highest request rates by 25.4–60.0% and 60.0–79.1% over per-input and per-workflow optimization, respectively. These results demonstrate that Aragog scales more effectively under increasing load.
@@
We observe several important trends. First, compared to per-input optimization, Aragog’s capacity improvements grow with workflow complexity. The throughput improvement increases from 50.0–50.1% for self-refine workflows to 54.0–76.3% for more complex workflows with either longer sequential chains (text-to-sql) or parallel branches (task decomposition and voting). This occurs because workflows with more stages exacerbate the limitations of per-input optimization’s static, upfront configuration. The configuration staleness increases with workflow complexity, leading to increasingly suboptimal configuration decisions. In contrast, Aragog exploits runtime reconfiguration and flexible request scheduling to maximize resource utilization across workflow stages, yielding higher throughput gains. Second, compared to per-workflow optimization, Aragog’s capacity improvements are more pronounced for simpler workflows (100.2–217.0%) than for complex workflows (78.1–128.6%). This is because complex workflows offer fewer opportunities for individual requests to exploit cheaper configurations, as tasks are harder. Similarly, the performance gap between per-workflow and per-input optimization narrows as workflow complexity increases.
@@
Latency reduction. Figure 12 presents detailed latency distributions across five workloads under the same request rates used in Figure 11. At low request rates where systems are underutilized, all approaches achieve comparable latency, as expected. However, as load increases, Aragog’s latency reductions across P25, median, and P95 all become increasingly pronounced. Under the highest load across all workloads, Aragog reduces P25, median, and P95 end-to-end latency by 23.6–54.6%, 32.5–58.8%, and 46.8–66.1% over per-input optimization and 58.4–75.1%, 60.0–78.9%, and 63.2–81.5% over per-workflow optimization, respectively. The latency gap widens most dramatically for the P95 metric, where Aragog achieves up to 66.1% and 81.5% reduction over the two baselines. This is because through runtime reconfiguration, Aragog avoids excessive queuing at individual models while maximizing resource utilization, effectively reducing tail latency compared to static baselines.
@@
Accuracy preservation. Figure 13 shows that Aragog maintains accuracy across all five workloads while achieving significant throughput and latency improvements. Aragog’s accuracy matches the most expensive configuration, with differences of at most 1 percentage across all workloads. This demonstrates that Aragog’s configuration predictor maintains high prediction accuracy while providing configuration flexibility for runtime scheduling.
@@ 5.3 Microbenchmarks
While the end-to-end evaluation demonstrates Aragog’s overall effectiveness, we conduct microbenchmarks to evaluate Aragog under more diverse setups, examining its performance across varying resource allocations, workload characteristics, and system parameters. We present results using the self-refine(HumanEval) and voting(MMLU-Pro) workflows as representative examples; the observed trends hold across all our workloads. Unless otherwise specified, all other experimental settings remain the same as described in §5.1.
System overheads. Aragog’s runtime overhead primarily stems from router inference Figure 14 shows the ratio of average routing latency to total end-to-end latency across varying request rates. Routing overhead remains minimal, accounting for at most 3.5% of total latency at low request rates and decreasing to below 1% at high request rates. Importantly, these percentages reflect the absolute routing latency; much of this overhead is overlapped with each request’s queueing delays. The ratio decreases as the request rate increases because queueing delays grow faster with load compared with routing latency, as routing computation is minimal compared to LLM inference in a workflow. These results demonstrate Aragog’s efficient routing optimizations.
@@
Beam search with different beam sizes. Figure 15 shows the impact of beam size on Aragog’s serving capacity improvement over two baselines. For self-refine on HumanEval, throughput improvement over per-input optimization grows from 16.3% to 51.2% as beam size increases from 1 to 8, while improvements over per-workflow optimization increase from 82.2% to 217.1%. Similarly, for voting on MMLU-Pro, improvements rise from 19.2% to 61.7% over per-input and 73.4% to 130.4% over per-workflow optimization. Performance gains plateau with increasing beam size, as beam search finds better configurations than greedy search (beam size 1) but with diminishing returns.
Figure 15: Impact of beam size on serving capacity improvement.
Mixture of Model Family. Figure 16 shows Aragog’s generality to heterogeneous model families by using models from mixed model families: Llama3 8B, Phi4 14B, and Qwen2.5 32B. While achieving the comparable accuracy as the most expensive configuration (within 1% loss), Aragog improves serving throughput by 33.7–68.5% over per-input optimization and 109.5–111.4% over per-workflow optimization. Additionally, Aragog reduces average latency under the highest request rates by 37.9–45.4% and 63.1–63.6% over per-input and per-workflow optimization, respectively. This demonstrates that Aragog generalizes effectively across diverse model architectures.
@@
Table 1: Sensitivity analysis of router’s embedding model, initial resource allocation, and configuration space size
(a) Impact of router architecture
Throughput $\uparrow$ Avg Latency $\downarrow$
@@
Arctic-Embed-L-v2.0 50.0% 217.0% 60.0% 79.1%
EmbeddingGemma 50.3% 216.8% 60.2% 79.5%
(b) Impact of GPU allocation scheme
Throughput $\uparrow$ Avg Latency $\downarrow$
@@
8 H100 GPUs 50.0% 217.0% 60.0% 79.1%
4 H100 GPUs 28.3% 205.2% 41.2% 78.6%
(c) Impact of configuration space size
Throughput $\uparrow$ Avg Latency $\downarrow$
@@
Full Space 50.0% 217.0% 60.0% 79.1%
Constrained Space 34.2% 164.6% 43.2% 51.7%
All improvements measured at the highest request rate from the end-to-end evaluation (§5.2). PI = per-input optimization, PW = per-workflow optimization. Crucially, all experiments maintain the same or better accuracy as the main evaluation.
Impact of router architectures. Table 1(a) compares two embedding models for the router: Arctic-Embed-L-v2.0 (our default) and Google’s EmbeddingGemma [embeddinggemma]. Both models have similar parameter counts and deliver comparable performance. The negligible performance differences demonstrate that Aragog’s routing mechanism is robust to the choice of embedding model, provided the embedding model can capture semantic similarity between workflow requests.
Impact of different GPU allocation schemes. We evaluate Aragog’s robustness to resource constraints by reducing GPU allocation from 8 to 4 H100 GPUs. With fewer GPUs, we reduce tensor parallelism accordingly: Qwen 32B from TP=4 to TP=2, Qwen 14B from TP=2 to TP=1, while Qwen 7B remains on a single GPU. This scenario favors the per-input baseline, which prioritizes smaller models based on static costs, while Aragog’s advantage comes from runtime reconfiguration that shifts load from overloaded small models to underutilized larger models. With 4 GPUs, limited capacity for larger models restricts such reconfiguration opportunities. Table 1(b) shows that Aragog still achieves significant improvements over per-input: 28.3% throughput gain and 41.2% latency reduction with 4 GPUs, compared to 50.0% and 60.0% with 8 GPUs. Gains over per-workflow optimization remain stable (205.2% vs. 217.0% in throughput, 78.6% vs. 79.1% in latency), as it often selects a fixed configuration with large models to meet worst-case accuracy requirements, making it similarly resource-constrained.
Impact of the size of configuration space. Aragog by default explores the full configuration space given a workflow and model candidates. Table 1(c) investigates performance under a constrained space with additional restrictions: for self-refine workflows, the critic and refine agents must use the same model; for voting workflows, all voters must use the same model. With the constrained space, Aragog achieves 34.2% and 43.2% improvements (throughput and latency) over per-input optimization versus 50.0% and 60.0% with the full space, and 164.6% and 51.7% over per-workflow optimization versus 217.0% and 79.1% with the full space. Despite reduced configuration flexibility, Aragog maintains significant advantages through joint scheduling that leverages each request’s available flexibility. This also demonstrates that Aragog’s joint scheduling can accommodate varying levels of configuration flexibility, offering a new dimension (configuration space size) for optimizing larger workflows.

### v2

@@ 4 Implementation
Figure 11: Aragog’s DSPy wrapper abstraction for enabling graph structure extraction and stage-wise scheduling.
We implement Aragog as an orchestration layer in 6.8K lines of Python and Triton code that sits between agentic workflow frontends and LLM serving engines, without modifying either.
Frontend and Workflow Specification. As shown in Figure 11, Aragog converts any DSPy program into its internal workflow representation, enabling runtime model reconfiguration and asynchronous stage-wise execution. We choose DSPy for its lightweight design and ease of use, making it popular for building agentic workflows. However. Aragog can support other frameworks by implementing framework-specific workflow graph extraction.
Router Inference. We implement dynamic batching for shared embedding model serving. To accelerate router classifier inference, we develop custom Triton kernels that fuse classifiers layer by layer. We leverage CUDA graphs to further reduce the kernel launching overheads.
Runtime Scheduler. The scheduler runs asynchronously with workflow execution. It maintains: (1) FIFO queues with look-ahead scheduling, (2) model engine states tracking occupancy, (3) configuration caches mapping requests to viable sets, and (4) exponential moving averages of queue metrics. Beam search uses priority queues, where each state encodes partial assignments as bit vectors for efficient manipulation.
Serving Backend Integration. Aragog uses SGLang [sglang], a widely used LLM serving engine as our serving backend. However, Aragog supports any serving backends (vLLM [vllm], TensorRT-LLM [tensorrt-llm] and HuggingFace TGI [TGI]) that provide OpenAI-compatible APIs needed for agentic workflow frontends.
@@ 5 Evaluation
We evaluated Aragog across a wide range of agentic workloads, model families, and varying request rates. Our key findings are:
@@
    Aragog improves maximum serving throughput by 42.8–76.3% and 78.1–217.0% over per-input and per-workflow optimizations, respectively.
@@
    Aragog maintains lower latency under increasing load. Under the highest load for all workloads, Aragog reduces P25, median, and P95 end-to-end latency by 23.6–60.1%, 32.5–71.1%, and 46.2–76.2% over per-input optimization and 58.4–82.8%, 60.0–86.1%, and 63.2–89.0% over per-workflow optimization, respectively.
@@
    Aragog preserves accuracy compared to always using the largest model, with maximum degradation of 2%.
@@
    Aragog demonstrates robust improvements over baselines across mixed model families, router architectures, configuration space sizes, and GPU allocation schemes.
@@ 5.1 Experimental Setup
Models. We use the Qwen-2.5 (7B, 14B, 32B) [qwen] and Llama 3 (3B, 8B, 70B) [llama3] model families. This provides a diverse range of capability-cost tradeoffs for each agent in the workflows. We use Snowflake Arctic-Embed-L-v2.0 [snowflake-model] as the router’s embedding model, which has a small number of parameters while providing high-quality embeddings and supporting relatively long context lengths. The router’s classifier is an MLP with 6 hidden layers that maps embeddings to configuration predictions. Additionally, we explore heterogeneous model options across different families, and different embedding models in §5.3.
Testbed. We conduct our experiments on an H100 GPU node, which contains 8 NVIDIA H100 80GB GPUs interconnected via NVLink and a 96-core Intel Sapphire Rapids CPU. We use SGLang as our model serving engine. For models larger than 7B, we enable tensor parallelism with a degree of 2 for 14B models and 4 for 32B and larger models, following standard practice. We dedicate the remaining GPU to router inference. For baselines without routing components, this GPU instead serves an additional 7B model to ensure fair resource allocation. We also explore different initial GPU allocation configurations in §5.3. As in prior work [vllm, sglang, autellix], we simulate request arrivals using a Poisson process with varying arrival rates to capture different load conditions.
Workflows and datasets. We evaluate Aragog across diverse workflow patterns (Figure 1) and datasets.
@@
    Self-refine for coding on HumanEval [humaneval] (164 programming problems), using a three-agent workflow with generation, critique, and refinement agents.
@@
    Text-to-SQL on Bird-SQL [bird] (780 samples), where query generation is decomposed into keyword extraction, column selection, SQL generation, and refinement stages.
@@
    Voting (four voters and one aggregator) on MMLU-Pro [mmlu-pro] (1500 sampled questions), where multiple reasoning paths are generated in parallel and aggregated.
@@
    Task decomposition on StrategyQA [strategyQA] (1500 sampled questions), where problems first pass through the decomposer, then invoke sub-task solvers in parallel, and finally the answer aggregator.
Baselines. We compare Aragog against two baselines that represent optimal performance for each category: (1) Per-workflow optimization: selects the cheapest configuration among all configurations that match the most expensive configuration’s accuracy, the configuration stays fixed for all requests; and (2) Per-input optimization: performs router-based selection for each request. Both baselines are given an oracle with perfect accuracy knowledge for selecting their best configurations, representing the performance upper bound of each approach. Additionally, we remove the router inference overhead for the per-input baseline. Moreover, all baselines incorporate the same state-of-the-art graph-aware scheduling for agentic workflows as Aragog (§3.2).
Figure 12: Average latency and throughput comparison under varying request rates and across five workflow dataset pairs and two model families. Each dot represents the average latency and throughput under a specific request rate.
Figure 13: Median latency for workloads and model families at varying loads. Error bars show P25 and P95.
@@ 5.2 End-to-End Performance
Figures 1214 compare Aragog with two baselines across five agentic workloads and two model families. Overall, Aragog significantly improves the maximal serving capacity and lowers latencies, while matching the accuracy when always using the largest model.
@@
Serving capacity improvement. Figure 12 shows throughput-latency curves under varying request rates across workloads and model families. For each workload, we sweep the request rate from low to high and measure the achieved throughput and average latency at each rate, shown as individual points on the curves. Across both Llama and Qwen model families, Aragog consistently achieves the highest throughput and lowest average latency under varying request rates. Specifically, Aragog improves maximum serving throughput by 42.8–76.3% over per-input optimization and 78.1–217.0% over per-workflow optimization. Additionally, Aragog reduces average latency under the highest request rates by 25.4–69.4% and 60.0–85.3% over per-input and per-workflow optimization, respectively. These results show that Aragog both scales efficiently under high load and its optimization generalizes across agentic workloads and model families.
@@
We observe several important trends across both model families. First, compared to per-input optimization, Aragog’s capacity improvements grow with workflow complexity. The throughput improvement increases from 42.8–50.1% for self-refine workflows to 54.0–76.3% for more complex workflows with longer sequential chains (text-to-sql) or parallel branches (task decomposition and voting). This is because workflows with more stages exacerbate the limitations of per-input optimization’s static, upfront configuration. Configuration staleness increases with workflow complexity, leading to increasingly suboptimal configuration decisions. In contrast, Aragog exploits runtime reconfiguration and flexible request scheduling to maximize resource utilization across stages, yielding higher throughput gains. Second, compared to per-workflow, Aragog’s capacity improvements are more pronounced for simpler workflows (100.2–217.0%) than complex ones (78.1–142.4%). This is because complex workflows offer fewer opportunities for requests to exploit cheaper configurations, as tasks are harder. Similarly, the gap between per-workflow and per-input optimization narrows as workflow complexity increases.
@@
Latency reduction. Figure 13 presents latency distributions across workloads and model families under the same request rates from Figure 12. At low request rates where systems are underutilized, all approaches achieve comparable latency, as expected. However, as load increases, Aragog’s latency reductions become increasingly pronounced. Under the highest load across all workloads, Aragog reduces P25, median, and P95 end-to-end latency by 23.6–60.1%, 32.5–71.1%, and 46.2–76.2% over per-input optimization and 58.4–82.8%, 60.0–86.1%, and 63.2–89.0% over per-workflow optimization. The improvements are most pronounced at P95, where Aragog achieves up to 89.0% reductions. This is because through flexible runtime reconfiguration, Aragog avoids excessive queuing at individual models, reducing tail latency compared to static baselines.
@@
Accuracy preservation. Figure 14 shows that Aragog maintains accuracy across all five workloads and two model families while achieving significant throughput and latency improvements. Aragog’s accuracy matches the most expensive configuration, with differences of at most 2%. This demonstrates that Aragog’s configuration predictor maintains high prediction accuracy while providing configuration flexibility for runtime scheduling.
@@ 5.3 Microbenchmarks
Results here use two representative workflows: self-refine on HumanEval and voting on MMLU-Pro, using models from the Qwen family; the observed trends hold for all our workloads. Unless otherwise specified, all other experimental settings remain the same as described in §5.1.
Routing overheads. Figure 15 shows the ratio of average routing latency to total end-to-end latency for different request rates. Routing overhead remains minimal, accounting for at most 3.5% of total latency at low request rates and <<1% at high request rates (when queueing delays are high). However, these percentages reflect absolute overheads; with Aragog, routing does not impose any additional delay on end-to-end latencies for requests. Indeed, at the highest request rates, routing completes 100% of the time during queuing of each request’s first stage; at the lowest request rates, Aragog’s budget-based routing (§3.1) does rarely terminate routing early, but even in such cases, 90% (on average) of routing work completes to ensure sufficient flexibility.
@@
Beam search with different beam sizes. Figure 16 shows the impact of beam size on Aragog’s wins. For self-refine on HumanEval, throughput improvement over per-input optimization grows from 16.3% to 51.2% as beam size increases from 1 to 8, while improvements over per-workflow optimization increase from 82.2% to 217.1%. Similarly, for voting on MMLU-Pro, gains rise from 19.2% to 61.7% over per-input and 73.4% to 130.4% over per-workflow. Performance gains plateau with increasing beam size, as beam search finds better configuration assignments than greedy search (beam size 1) but with diminishing returns.
Figure 16: Impact of beam size on throughput improvement.
Mixture of Model Family. Figure 17 shows Aragog’s generality to heterogeneous model families by using models from mixed model families: Llama3 8B, Phi4 14B, and Qwen2.5 32B. While maintaining the accuracy as the most expensive configuration, Aragog improves serving throughput by 33.7–68.5% over per-input baseline and 109.5–111.4% over per-workflow baseline. Additionally, Aragog reduces average latency under the highest request rates by 37.9–45.4% and 63.1–63.6% over per-input and per-workflow optimization, respectively.
@@
Table 1: Throughput improvement under different router embedding models, GPU allocations, and configuration space sizes
(a) Impact of router embedding models
Self-refine Voting
@@
Arctic-Embed-L-v2.0 50.0% 217.0% 60.0% 128.6%
EmbeddingGemma 50.3% 216.8% 58.3% 126.7%
(b) Impact of GPU allocation schemes
Self-refine Voting
@@
8 H100 GPUs 50.0% 217.0% 60.0% 128.6%
4 H100 GPUs 28.3% 205.2% 32.1% 117.4%
(c) Impact of configuration space sizes
Self-refine Voting
@@
Full Space 50.0% 217.0% 60.0% 128.6%
Constrained Space 34.2% 164.6% 52.1% 106.5%
All improvements measured at the highest request rate from the main evaluation (§5.2). PI = per-input baseline, PW = per-workflow baseline. All experiments maintain the same or better accuracy as the main evaluation.
Impact of router architectures. Table 1(a) compares two embedding models for the router: Arctic-Embed-L-v2.0 (our default) and Google’s EmbeddingGemma [embeddinggemma]. Both have similar parameter counts and deliver comparable throughput performance (within 1.9%). The negligible performance differences show that Aragog’s routing mechanism is robust to the choice of embedding model, provided that it can capture semantic similarity between workflow requests.
Impact of different GPU allocation schemes. We evaluate Aragog’s robustness to resource constraints by reducing GPU allocation from 8 to 4 H100 GPUs. With fewer GPUs, we reduce tensor parallelism accordingly: Qwen 32B from TP=4 to TP=2, Qwen 14B from TP=2 to TP=1, while Qwen 7B remains on a single GPU. This scenario favors the per-input baseline, which prioritizes smaller models based on static costs, while Aragog’s advantage comes from runtime reconfiguration that shifts load from overloaded small models to underutilized larger models. With 4 GPUs, limited capacity for larger models restricts such reconfiguration opportunities. Table 1(b) shows that Aragog still achieves significant throughput improvements over the per-input baseline with 4 GPUs: 28.3% for self-refine and 32.1% for voting, compared to 50.0% and 60.0% with 8 GPUs. Gains over per-workflow optimization remain stable across both workloads (205.2% and 117.4% with 4 GPUs vs. 217.0% and 128.6% with 8 GPUs), as it selects a fixed configuration with large models to meet worst-case accuracy requirements, making it similarly resource-constrained.
Impact of the size of configuration space. Aragog by default explores the full configuration space given a workflow and model candidates. Table 1(c) investigates performance under a constrained space with additional restrictions: for self-refine workflows, the critic and refine agents must use the same model; for voting workflows, all voters must use the same model. With the constrained space, Aragog achieves 34.2% and 52.1% throughput improvements over per-input optimization for self-refine and voting, respectively, versus 50.0% and 60.0% with the full space. Gains over per-workflow optimization decrease from 217.0% and 128.6% to 164.6% and 106.5% for the two workloads. Despite reduced configuration flexibility, Aragog maintains significant advantages through joint scheduling that leverages each request’s available flexibility. This also demonstrates that Aragog allows users to configure the configuration space size, which can be useful for managing complexity in larger workflows with more agents.